Ticket Number: HEA-NOC/20080201-3 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver: HEA-NOC
Ticket Opened: 20080201 12:24 UTC Problem Start: 20080131 18:04 UTC
Ticket Update: 20080201 12:45 UTC Problem End: 20080131 19:32 UTC
Site/line: IT Cork, LIT, UCC, UL, UL Mary Immaculate, UL Outreach
During the problem period there were 7 flaps on this link, each lasting
less than 10 seconds.
Loss of resilience to the affected sites listed above, and their clents.
Traffic will have re-routed over alternative links.
20080201 11:29 LT The affected link is carried over fibre from e-net and
ESBT. Investigating the cause.
20080201 12:43 LT Seeking assistance from NSN on interpreting ADVA alarms,
call reference HEANET-J-004-2008.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20080201-3
HEAnet Limited. Registered in Ireland, No. 275301.